OpenStudy (rational):

show that \((p-2)!\) is divisible by \(p-1\) for all primes \(p\gt 5\)

OpenStudy (ikram002p):

same as showing that (p-1)! is divisible by p hehe ok trying to restore Wilson proof

OpenStudy (rational):

(p-1)! is never divisible by p by wilson thm because (p-1)! = -1 (mod p)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} (p-2)! \pmod {(p-1)}\hspace{.33em}\\~\\ \implies (p-2)(p-1)p! \pmod {(p-1)}\hspace{.33em}\\~\\ \implies (p-2)(p-1)p! \pmod {(p-1)}\equiv 0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (tkhunny):

p = 3 p-2 = 1 1! = 1 1 is not divisible by 2 Is there an appropriate minimum p?

OpenStudy (rational):

yes sorry p > 5 @tkhunny

OpenStudy (rational):

let me update the question

OpenStudy (ikram002p):

oh hehe misunderstood

OpenStudy (rational):

also mathmath's method wont work, we can't multiply a number..

OpenStudy (ikram002p):

ok (p-2)! =1*2*3*.....*(p-3)*(p-2) humm (thinking)

OpenStudy (rational):

6 is not prime

OpenStudy (ikram002p):

p is prime :\

OpenStudy (rational):

updated the question again sry ;p

OpenStudy (mathmath333):

ok u want primes

OpenStudy (ikram002p):

ok got it p-1 is even composite thus p-1=2*n such that n<p-2 hence (p-2 )! |n and (p-2)!|2 since p>5 there would be extra 2's in 4 so it wont matter if n is even or not

OpenStudy (ikram002p):

wew! so cute

OpenStudy (rational):

can i have a bit more rigorous proof

OpenStudy (ikram002p):

ok p is odd thus p-1 is even p-1=2*n (such that n <p-2) so \( n \in (1,2,3,.....,p-3) \\ n| (1*2*3*.....*p-2)=(p-2)!\) now when n is odd n>2 thus n can be any number btw 3 and p-2 (so n wont be 2) hence 2*n|(p-2)! when n is even p-1>=8 thus n btw 4 and p-2 (so n wont be 2) hence 2*n|(p-2)!

OpenStudy (ikram002p):

hmm :\ sounds clear to me idk what ur looking for

OpenStudy (rational):

Okay nice that will do

OpenStudy (ikram002p):

hmm

OpenStudy (ikram002p):

do u have some other idea ?